Use Warshall’s algorithm to find the transitive closure of the relation R = {(1,1),(1,4), (2,5),(2,3),(3,1),(3,5),(3,4),(4,2)} on the set A = {1,2,3,4,5}.
1
Expert's answer
2020-06-16T17:37:25-0400

"A=\\\{1,2,3,4,5\}\\
R=\{(1,1),(1,4), (2,5),(2,3),(3,1),(3,5),\\
(3,4),(4,2)\}"

"w_{ij}^{(k)}=\left\{\begin{matrix}
 w_{ik}^{(k-1)}\lor w_{kj}^{(k-1)}, \quad if \quad w_{kj}^{(k-1)}=1\\
 w_{ij}^{(k-1)}, \quad if \quad w_{ik}^{(k-1)}=0
\end{matrix}\right\}"



"W_0=\begin{matrix}
 &1_{-}&2_{-}&3_{-}&4_{-}&5_{-} \\
 1_{-} & 1&0&0&1&0\\
2_{-}&0&0&1&0&1\\
3_{-}&1&0&0&1&1\\
4_{-}&0&1&0&0&0\\
5_{-}&0&0&0&0&0
\end{matrix}"

Find "W_1"

"W_1=\begin{matrix}
 &1_{-}&2_{-}&3_{-}&4_{-}&5_{-} \\
 1_{-} & 1&0&0&1&0\\
2_{-}&0&0&1&0&1\\
3_{-}&1&0&0&1&1\\
4_{-}&0&1&0&0&0\\
5_{-}&0&0&0&0&0
\end{matrix}"


"W_2=\begin{matrix}
 &1_{-}&2_{-}&3_{-}&4_{-}&5_{-} \\
 1_{-} & 1&0&0&1&0\\
2_{-}&0&0&1&0&1\\
3_{-}&1&0&0&1&1\\
4_{-}&0&1&1&0&1\\
5_{-}&0&0&0&0&0
\end{matrix}"


"W_3=\begin{matrix}
 &1_{-}&2_{-}&3_{-}&4_{-}&5_{-} \\
 1_{-} & 1&0&0&1&0\\
2_{-}&1&0&1&1&1\\
3_{-}&1&0&0&1&1\\
4_{-}&1&1&1&1&1\\
5_{-}&0&0&0&0&0
\end{matrix}"


"W_4=\begin{matrix}
 &1_{-}&2_{-}&3_{-}&4_{-}&5_{-} \\
 1_{-} & 1&1&1&1&1\\
2_{-}&1&1&1&1&1\\
3_{-}&1&1&1&1&1\\
4_{-}&1&1&1&1&1\\
5_{-}&0&0&0&0&0
\end{matrix}"


"W_4=W_5"